The table shows the probability distribution for the number of heads when four coins are tossed. What is the probability that no more than two heads show on a random toss?
X = Number of Heads: 0, 1, 2, 3, 4
Probability: 0.0625, 0.25, 0.375, 0.25, 0.0625

My Work: My answer was 0.3125 because worked it out like this
P(X<2) = P(X=0) + P(X=1)
= 0.0625 + 0.25
= 0.3125
I just want to make sure that I have it right. Thank you for helping me.

Answer by Edwin McCravy(20056)

No, the problem states "no more that 2", not "less than 2".
That's "less than or equal 2".  So it's P(X ≤ 2), not P(X < 2).


P(X ≤ 2) = P(X=0) + P(X=1) + P(X=2)
= 0.0625 + 0.25 + 0.375
= 0.8675

Edwin